susan paid .49 per page for color copy and .9 per page for black and white copy. If susn bill for 90 copies was $12.90. How may copies of each type were made

Answer by macston(5194)

.
C=number of color copies; B=number of black and white copies
B+C=90
B=90-C Use this to substitute for B.
$0.09B+$0.49C=$12.90 Substitute for B.
$0.09(90-C)+$0.49C=$12.90
$8.10-$0.09C+$0.49C=$12.90 Subtract $8.10 from each side.
$0.40C=$4.80 Divide each side by $0.40
C=12 ANSWER 1: There were 12 color copies.
B=90-C=90-12=78 ANSWER 2: There were 78 black and white copies.
.
CHECK:
$0.09B+$0.49C=$12.90
$0.09(78)+$0.49(12)=$12.90
$7.02+$5.88=$12.90
$12.90=$12.90